On June 1, Vandervelde Corporation (a U.S.-based manufacturing firm) received an order to sell goods to a foreign customer at a price of 205,000 leks. Vandervelde will ship the goods and receive payment in three months on September 1. On June 1, Vandervelde purchased an option to sell 205,000 leks in three months at a strike price of $0.89. It properly designated the option as a fair value hedge of a foreign currency firm commitment. The fair value of the firm commitment is measured by referring to changes in the spot rate. Relevant exchange rates and option premiums for the lek are as follows:

Date Spot Rate Put Option Premium
for September 1
(strike price $0.89)
June 1 $ 0.89 $ 0.020
June 30 0.83 0.072
September 1 0.78 N/A


Vandervelde’s incremental borrowing rate is 12 percent. The present value factor for two months at an annual interest rate of 12 percent (1 percent per month) is 0.9803. Vandervelde Corporation must close its books and prepare its second-quarter financial statements on June 30.

  1. Prepare journal entries for the foreign currency option and firm commitment.

  2. What is the impact on net income over the two accounting periods?

  3. What is the net cash inflow resulting from the sale of goods to the foreign customer?
Answer #1

a.

For June 1, calculate the value of the foreign currency option:

Foreign Currency Option = 205,000 * 0.020 = $ 4,100

Journal entry for June 1:

Credit $ Date 1-Jun Account name Foreign Currency Option Cash Debit $ 4,100 4,100

There is no need to record the sales agreement since it is considered as an executive contract.

Since the spot rates decreased between the two dates of June, there is a loss on the firm commitment.

Loss on firm commitment = [ 205,000 * ( 0.83 - 0.89 ) ] * 0.9803 = ($12,057.70)

Since the Put Option Premium rates increased in value between the two dates in June, there is a gain on the foreign currency option.

Gain on foreign currency option = 205,000 * ( 0.072 - 0.020) = $10,660

Prepare the appropriate journal entry for June 30

Credit $ Date 30-Jun Account name Loss on Firm Commitment Firm Commitment Debit $ 12,057.70 12,058 10,660 Foreign Currency Op

For September 1, the spot rate values have decreased from the previous two dates, therefore, causing a loss on the value of the firm commitment.

Loss on firm commitment = [ 205,000 * ( 0.78 - 0.89) ] + 12,057.7 = ($10,492.30)

Since the value of the Put option has increased, there is a gain on the foreign currency option

Gain on Foreign Currency = [ 205,000 * ( 0.89 - 0.78 ) - ( 205,000 * 0.072) = $ 7,790

Calculate the amount to be recorded as sales on September 1:

Foreign Currency = 205,000 * 0.78 = $159,900

Calculate the amount of cash to be recorded on September 1:

Cash = 205,000 * $0.89 = $182,450

It is taken as the rate of $0.89 because that was the agreed upon strike price at the time of purchase and not at the spot rate of $0.78.

Calculate the amount that will be reported from the foreign currency option:

Foreign Currency Option = Cash - Foreign Currency

= $182,450 - $159,900

= $22,550

This amount is also considered as the adjustment to the net income.

Prepare an appropriate Journal entry detailing the transactions:

Credit $ Date 1-Sep Account name Loss on Firm Commitment Firm Commitment Debit $ 10,492.30 10,492.30 7,790 Foreign Currency O

.

Calculate the impact of net income over the second quarter:

Particulars Loss on Firm Commitment Gain on Foreign Currency Option Impact on Net Income Amount $ -12,057.70 10,660 -1,397.70

The impact on net income for the second quarter is a negative $1,397.7

Calculate the impact of net income over the third quarter:

Particulars Sales Loss on Firm Commitment Gain on Foreign Currency Option Adjusted to Net Income Impact on Net Income Amount

The impact of net income over the third quarter is $179,747.70 an amount that more than makes up for the loss on net income for the second quarter.

.

Calculate the net cash inflow from the sale of goods to the foreign customer.

Net cash inflow = 182,450 - 4,100 = $178,350.

*** End ***

Do leave a thumbs up : )